# Exploiting Unlabeled Data with Multiple Expert Teachers for Open Vocabulary Aerial Object Detection and Its Orientation Adaptation

### v1

@@ III Method
We start by presenting the problem setting in Section IV-A). Next, we provide an overview of our CastDet framework in Section IV-B, which is composed of several key components: i) the student, a two-stage open-vocabulary detector, discussed in Section IV-B1; ii) the localization teacher, as detailed in Section IV-B2; and iii) the dynamic pseudo-label queue for maintaining high-confidence pseudo-labels, outlined in Section IV-B3. Finally, we elaborate on two open-vocabulary aerial object detection tasks: horizontal open-vocabulary detection in Section IV-C and oriented open-vocabulary detection in Section IV-D.
IV Methodology
IV-A Preliminaries
@@
Given an image $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$ , our objective is to train a detector capable of detecting both base ($\mathcal{C}_{\text{base}}$) and novel ($\mathcal{C}_{\text{novel}}$) objects within the image. This task involves solving two subproblems: i) localize all objects with precise bounding boxes, where each box is represented by coordinates $b_{i}\in\mathbb{R}^{4}$ for horizontal objects, and $b_{i}\in\mathbb{R}^{5}$ for oriented objects; and ii) classify the $i$-th object with the category label $c_{i}\in\mathbb{R}^{\mathcal{C}_{\text{test}}}$, where $\mathcal{C}_{\text{test}}=\mathcal{C}_{\text{base}}\cup\mathcal{C}_{\text{novel}}$ and $\mathcal{C}_{\text{base}}\cap\mathcal{C}_{\text{novel}}=\emptyset$. During training, we utilize a labeled dataset =$\mathcal{L}=\{(\mathbf{x},\{(b,c)_{k}\})_{i}\}_{i=1}^{|\mathcal{L}|}$ on a set of base categories, i.e., $c_{k}\in\mathbb{R}^{\mathcal{C}_{\text{base}}}$. Additionally, we incorporate an unlabeled dataset $\mathcal{U}=\{\mathbf{x}_{i}\}_{i=1}^{|\mathcal{U}|}$, which consists of images without bounding box annotations. The combined training dataset is represented as $\mathcal{D}_{\text{train}}=\mathcal{L}\ \cup\ \mathcal{U}$.
IV-B CLIP-Activated Student Teacher Framework
@@
Figure 3: Overall architecture of CastDet. In each training iteration, the data batch consists of three data flow: labeled data with annotations, unlabeled data, and data sampled from the dynamic label queue. The labeled images are directly used for the student network training ($\mathcal{L}_{s}$), while two sets of pseudo-labels of unlabeled data are predicted through the localization teacher and external teacher. One supervises the student ($\mathcal{L}_{u}$), and the other is pushed into the dynamic label queue. Simultaneously, samples are randomly selected from the dynamic label queue to enhance the student’s ability to detect novel objects ($\mathcal{L}_{d}$).
@@
Architecture Overview. Fig. 3 illustrates an overview of our CastDet framework, which consists of a student model and two teacher models: a localization teacher and an external teacher. The student model is a two-stage object detection model (e.g., Faster R-CNN or Oriented RCNN) with a modified class-agnostic bounding box regression head ($\Phi_{\text{LOC}}$) and a semantic classifier ($\Phi_{\text{CLS}}$). It is trained on both labeled and unlabeled samples with pseudo boxes and category labels generated by the localization teacher and external teacher. The localization teacher is an exponential moving average (EMA) of the student model [70], which can aggregate the history information during the training iterations to obtain better and stable representations, ensuring the quality of the pseudo-labels. During training, the localization teacher generates two sets of pseudo-labels for the unlabeled images, one for training the student model and another with pseudo boxes serve as input to the external teacher for the pseudo category label generation. The external teacher is a frozen RemoteCLIP model [33], a vision-language fundamental model pre-trained on large-scale remote sensing image-text pairs, following the CLIP framework. It provides strong open-vocabulary classification ability by comparing the similarity between image and category embeddings. Additionally, a dynamic label queue is used to store the pseudo-labels generated by the external teacher, facilitating to maintain high-quality pseudo-labels and transfer data samples for the student model training.
IV-B1 Student
@@
The student model is a two-stage object detector, such as Faster R-CNN and Oriented-RCNN, comprising a backbone encoder, a region proposal network (RPN), and a box head module:
$\{\hat{y}_{1},\ldots,\hat{y}_{n}\}=\Phi_{\text{BOX}}\circ\Phi_{\text{RPN}}\circ\Phi_{\text{ENC}}(\mathbf{x})$
@@
To construct an open-vocabulary detector, we incorporate two key components into the box head: (1) a class-agnostic box regression head ($\Phi_{\text{LOC}}$) to generate class-agnostic box coordinates, and (2) a semantic classifier head ($\Phi_{\text{CLS}}$) to classify these objects beyond seen categories, i.e., open-vocabulary classification.
@@
Class-agnostic box regression head ($\Phi_{\text{LOC}}$) takes the RoI (Region of Interest) features as input and produces refined box coordinates for each object. To handle a variable number of object categories, $\Phi_{\text{LOC}}$ shares parameters across all categories. Specifically, the regression box is represented as $b_{i}\in\mathbb{R}^{k}$ instead of $b_{i}\in\mathbb{R}^{k|\mathcal{C}_{\text{test}}|}$ for each box $i$, where $k\in\{4,5\}$ is the dimension of the box. As described in [84, 24], this approach simplifies the model and enhances its versatility.
@@
Semantic classifier head ($\Phi_{\text{CLS}}$) is responsible for classifying RoI regions beyond a predefined set of categories. Following Detic[24], we employ semantic embeddings for category vocabularies as the weight of the last fully connected layer, allowing flexible category expansion. The semantic embeddings are generated by two steps: (1) filling the concept with a predefined prompt template “a photo of [category]”, and (2) encoding the text descriptions into semantic embeddings using the text encoder of RemoteCLIP.
@@
Given a set of RoI visual features $\{v_{i}\}_{i=1}^{k}$, the similarity score is calculated as
where $\tau$ is the temperature parameter that controls the range of the logits, optimized during training as a log parameterized multiplicative scalar as in [82].
IV-B2 Localization Teacher
The quality region proposals is fundamental for effectively detecting novel objects, as it directly impacts the subsequent box refinement and classification. However, as shown in Fig. 2(c), the RPN recall of novel categories in aerial images is significantly lower than that in natural images. To address this problem, we employ a robust teacher to generate better pseudo-boxes for the student training. Our goal, then, is to derive a better teacher model from the student model without introducing additional training costs. A naive solution is to employ a frozen teacher model trained on base categories, but this is insufficient for open-vocabulary detection due to its lack of flexibility in learning novel categories. Thus, we opt for a continuously update mechanism, enabling the teacher to evolve and adapt to novel concepts, thereby enhancing its effectiveness in a broader detection context.
Exponential Moving Average. Inspired from [70, 90], we employ a momentum update strategy for the teacher model, that is, the teacher is updated by an exponential moving average of the student during training. Formally, denoting the parameters of the teacher model as $\theta^{\prime}$ and those of the student model as $\theta$. Only the parameters $\theta$ are updated by back-propagation, and $\theta^{\prime}$ as a weighted average of successive weights of $\theta$ at each training iteration $t$:
$\theta_{t}^{\prime}=\alpha\theta_{t-1}^{\prime}+(1-\alpha)\theta_{t}$ (3)
Here $\alpha\in[0,1)$ is a momentum coefficient.
The momentum update described in Equ. (3) enhances the teacher model in several significant ways compared to a frozen teacher. 1) it enables the teacher to leverage unlabeled data effectively, enhancing the performance of the model with fewer annotated images; 2) it accumulates historical information of the student model without additional training costs, thereby obtaining better intermediate representations and more robust predictions [70, 28]; and 3) it facilitates on-line learning, allowing the model to continuously adapt and scale to an increasing number of novel concepts.
Consistency Training. We apply consistency training [29, 30] that encourages the student to predict the same prediction as the teacher. Specifically, given an unlabeled image $\mathbf{x}$, it first go through weak augmentation and serve as input for the localization teacher to generate pseudo-labels. The pseudo-labels are then used to supervise the student model, which takes the same image with a strong augmentation as input. More formally, we define the prediction of the teacher model as $\mathbf{\hat{y}}=f(T^{\prime}(\mathbf{x}),\theta^{\prime})$. We then minimize the consistency cost $J$ between $\mathbf{\hat{y}}$ and the predictions of the student model, i.e.,
$J(\theta)=\mathcal{L}(f(T(\mathbf{x}),\theta),\mathbf{\hat{y}})$ (4)
where $T$ and $T^{\prime}$ denote strong and weak augmentation, respectively.
IV-B3 Dynamic pseudo-label Queue
@@
Figure 4: Workflow of the dynamic label queue. Step1: filter certain high-quality proposal boxes generated by the localization teacher, and employ RemoteCLIP to classify corresponding cropped images as pseudo-labels. Step2: dynamically update the pseudo-labels into the queue, and randomly sample a batch of pseudo-labels for the student training.
@@
The dynamic label queue is responsible for storing, dynamically updating, and transferring pseudo-labels. At each training iteration, the localization teacher and external teacher generate high-quality pseudo-labels for the unlabeled images, which are then pushed into the label queue. Following this, a random batch of pseudo-labels is sampled from the dynamic queue for the student training, as illustrated in Fig. 4.
@@
Generate pseudo-labels. The localization teacher take unlabeled images with weak augmentation $T^{\prime}(\mathbf{x})$ as input, where the backbone first extracts features, the RPN then generates a set of proposals $\{p_{i}\}$ and extract RoI features $\mathbf{f}^{P}=\Phi_{\text{RPN}}\circ\Phi_{\text{ENC}}(T^{\prime}(\mathbf{x}))$ of these proposals through RoI pooling. Subsequently, the class-agnostic regression branch predicts their refined coordinates $\{\hat{b}_{i}\}$. To improve the accuracy of pseudo-labels while reducing computational complexity, we use a box filter to select $k$ candidates and obtain a set of image crops $\{\mathbf{x}^{I}_{i}\}_{i=1}^{k}$ by cropping the corresponding regions of the image. The category labels for these regions are then predicted by the RemoteCLIP.
@@
For each cropped image $\mathbf{x}^{I}_{i}$, the visual feature $v_{i}$ is computed via the visual encoder of the RemoteCLIP, denoted as $v_{i}=\Phi_{\text{clip}}^{\text{visual}}(\mathbf{x}^{I}_{i})$. At the same time, text embeddings $t_{j}$ corresponding to category descriptions are generated, as detailed in Sec. IV-B. The prediction probability for each category $j$ given the cropped image $i$ is the softmax value for similarity between the visual feature and text embeddings:
@@
We further apply a threshold $p_{0}$ to filter pseudo-labels, which is set relatively high to ensure the accuracy of pseudo-labels,
$\hat{y}=\{(\hat{b}_{i},\hat{c}_{i})|\hat{p}_{i}\geq p_{0}\}_{i=1}^{\hat{k}},\ \hat{c}_{i}=\arg\max_{j}\hat{p}_{ij},\ \hat{p}_{i}=\max_{j}\hat{p}_{ij}$ (6)
where $\hat{p}_{i}$, $\hat{c}_{i}$ are the prediction score and category label, respectively. Finally, the image with its pseudo-labels $(\mathbf{x},\hat{y})$ is pushed into the dynamic label queue.
@@
This dynamic process allows the queue to progressively accumulate richer and more accurate pseudo-labels as the model iterates, with visualization results shown in Fig. 10. The continuous improvement benefits the entire system, enabling the student and localization teacher to learn from diverse, high-quality pseudo-labels.
IV-C Horizontal Open-vocabulary Detection
Horizontal object detection aims to detect all objects in a image and represent their location with horizontal bounding boxes, i.e., $b_{i}\in\mathbb{R}^{4}$. In theory, the detector can be any mainstream object detectors [26, 43, 41], here we use Faster R-CNN [26] as the base model to illustrate our method. To enable open-vocabulary detection, we incorporate external knowledge that may carry information about novel categories from unlabeled images into the training process. More specifically, the localization teacher first generates a set of box candidates, which are then filtered based on their confidence scores. Only those with high confidence are forwarded to the external teacher for category prediction. Subsequently, we leverage both unlabeled data with these high-quality pseudo-labels and labeled data to optimize the model via a hybrid training mechanism.
IV-C1 Horizontal Box Selection Strategies
The quality of pseudo-boxes is vital, as it directly affects the accuracy of the category labels generated by the external teacher, subsequently influencing the overall performance of the model. Given this context, we utilize several box confidence scores to filter high-quality pseudo boxes, including RPN, RJV and BJV Score, as shown in Fig. 5(a)$\sim$(c).

### v2

@@ 3 Methodology
We start by presenting the open-vocabulary object detection problem setting in Section 3.1). Next, we provide the overview of our CastDet framework in Section 3.2, which is composed of three key components: i) the student module, a two-stage open-vocabulary detector, introduced in Section 3.2.1; ii) the localization teacher model, as detailed in Section 3.2.2; and iii) the dynamic pseudo-label queue for maintaining high-confidence pseudo-labels, outlined in Section 3.2.3. Finally, we elaborate on two tasks: horizontal open-vocabulary detection in Section 3.3 and oriented open-vocabulary detection in Section 3.4.
3.1 Definition of Open Vocabulary Aerial Object Detection
@@
Open Vocabulary Aerial Object Detection (OVAD) aims to extend traditional aerial object detection within base classes to detect novel objects emerged in earth observation, addressing the unique challenges inherent to aerial scenarios, including weak feature appearance, orientation diversity, and the difficulty of obtaining sufficient labeled data. Given an image $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$ , the goal of the newly defined OVAD task is to train a detector capable of detecting both base ($\mathcal{C}_{\text{base}}$) and novel ($\mathcal{C}_{\text{novel}}$) objects within the image. This task involves solving two subproblems: i) localize all objects with precise bounding boxes, where each box is represented by coordinates $b_{i}\in\mathbb{R}^{4}$ for horizontal objects, and $b_{i}\in\mathbb{R}^{5}$ for oriented objects; and ii) classify the $i$-th object with the category label $c_{i}\in\mathbb{R}^{\mathcal{C}_{\text{test}}}$, where $\mathcal{C}_{\text{test}}=\mathcal{C}_{\text{base}}\cup\mathcal{C}_{\text{novel}}$ and $\mathcal{C}_{\text{base}}\cap\mathcal{C}_{\text{novel}}=\emptyset$. During training, we utilize a labeled dataset $\mathcal{L}=\{(\mathbf{x},\{(b,c)_{k}\})_{i}\}_{i=1}^{|\mathcal{L}|}$ on a set of base categories, i.e., $c_{k}\in\mathbb{R}^{\mathcal{C}_{\text{base}}}$. Additionally, we incorporate an unlabeled dataset $\mathcal{U}=\{\mathbf{x}_{i}\}_{i=1}^{|\mathcal{U}|}$, which consists of images without bounding box annotations. The combined training dataset is represented as $\mathcal{D}_{\text{train}}=\mathcal{L}\ \cup\ \mathcal{U}$.
3.2 CLIP-Activated Student Teacher Framework
@@
Figure 4: Overall architecture of CastDet. In each training iteration, the data batch consists of three data flow: labeled data with annotations, unlabeled data, and data sampled from the dynamic label queue. The labeled images are directly used for the student model training ($\mathcal{L}_{s}$), while two sets of pseudo-labels of unlabeled data are predicted through the localization teacher model and external teacher model. One supervises the student model ($\mathcal{L}_{u}$), and the other is pushed into the dynamic label queue. Simultaneously, samples are randomly selected from the dynamic label queue to enhance the student’s ability to detect novel objects ($\mathcal{L}_{d}$).
@@
Architecture Overview. Fig. 4 illustrates the overview of our CastDet framework, which consists of a student model and two teacher models: a localization teacher and an external teacher. The student model is a two-stage object detection model (e.g., Faster R-CNN or Oriented RCNN) with a modified class-agnostic bounding box regression head ($\Phi_{\text{LOC}}$) and a semantic classifier ($\Phi_{\text{CLS}}$). It is trained on both labeled and unlabeled samples with pseudo boxes and category labels generated by the localization teacher model and external teacher model. The localization teacher model is an exponential moving average (EMA) of the student model [tarvainen2017mean], which can aggregate the history information during the training iterations to obtain better and stable representations, ensuring the quality of the pseudo-labels. During training, the localization teacher model generates two sets of pseudo-labels for the unlabeled images, one is used for training the student model, and another serves as input to the external teacher model for generating pseudo category labels. The external teacher model is a frozen RemoteCLIP [liu2023remoteclip], a vision-language fundamental model pre-trained on large-scale remote sensing image-text pairs, following the CLIP framework. It provides strong open-vocabulary classification ability by comparing the similarity between image and category embeddings. Additionally, a dynamic label queue is used to store the pseudo-labels generated by the external teacher model, facilitating to maintain high-quality pseudo-labels and transfer data samples for the student model training.
3.2.1 Student
@@
The student model is a two-stage object detector, represented by Faster R-CNN or Oriented-RCNN, comprising a backbone encoder ($\Phi_{\text{ENC}}$), a region proposal network ($\Phi_{\text{RPN}}$), and a box head module ($\Phi_{\text{BOX}}$):
$\{\hat{y}_{1},\ldots,\hat{y}_{n}\}=\Phi_{\text{BOX}}\circ\Phi_{\text{RPN}}\circ\Phi_{\text{ENC}}(\mathbf{x}),$
where the operator $\circ$” denotes function composition, i.e., the input $\mathbf{x}$ is processed sequentially by $\Phi_{\text{ENC}}$, $\Phi_{\text{RPN}}$, and $\Phi_{\text{BOX}}$.
@@
To construct an open-vocabulary detector capable of detecting novel objects beyond the close-set categories, we incorporate two key components into the box head: i) a class-agnostic box regression head ($\Phi_{\text{LOC}}$) to generate class-agnostic box coordinates, and ii) a semantic classifier head ($\Phi_{\text{CLS}}$) to classify these objects beyond seen categories, i.e., open-vocabulary classification.
@@
Class-agnostic box regression head ($\Phi_{\text{LOC}}$) takes the RoI (Region of Interest) features as input and outputs refined box coordinates for each object. To handle a variable number of object categories, $\Phi_{\text{LOC}}$ shares parameters across all categories. Specifically, the regression box is represented as $b_{i}\in\mathbb{R}^{k}$ instead of $b_{i}\in\mathbb{R}^{k|\mathcal{C}_{\text{test}}|}$ for each box $i$, where $k\in\{4,5\}$ is the dimension of the box, with $k=4$ for horizontal boxes and $k=5$ for oriented boxes. This approach is also in line with the previous works [feng2022promptdet, zhou2022detic].
@@
Semantic classifier head ($\Phi_{\text{CLS}}$) is devised for classifying RoI regions beyond a predefined set of categories. Following Detic[zhou2022detic], we employ semantic embeddings for category vocabularies as the weights of the last fully connected layer for flexible category expansion. The semantic embeddings are generated by two steps: i) filling concept into the predefined prompt template “a photo of [category]”, and ii) extracting semantic embeddings of text descriptions using the text encoder of RemoteCLIP.
$t_{j}=\Phi_{\text{clip}}^{\text{text}}\circ\Phi_{\text{clip}}^{\text{tokenizer}}(\text{a photo of [category]}),$ (1)
where $\Phi_{\text{clip}}^{\text{tokenizer}}(\cdot)$ denotes the text tokenizer that maps text into tokens, $\Phi_{\text{clip}}^{\text{text}}(\cdot)$ is the text encoder that converts tokens into the final semantic embedding $t_{j}$ for category $j$.
Given a set of RoI visual features $\{v_{i}\}$, the similarity score between $i$-th RoI feature and $j$-th category is calculated as:
where $\tau$ is the temperature parameter that controls the range of the logits, and it is optimized as a log parameterized multiplicative scalar as in [CLIP_radford2021learning].
Figure 5: Augmentation strategies used in our approach.Weak augmentation applies only random flip (highlighted in orange box), whereas strong augmentation incorporates all augmentation techniques.
3.2.2 Localization Teacher
The quality of region proposals is significant for effectively detecting novel objects, as it directly impacts the subsequent box refinement and classification. However, as shown in Fig. 2, the RPN recall of novel categories in aerial images is significantly lower than that in natural images. To address this problem, we employ a robust teacher model to generate better pseudo-boxes for training the student model. Successively, deriving a better teacher model from the student one without introducing additional training costs is crucial. A naive solution is to employ a frozen teacher model trained on base categories, however, a issue rises that it lacks flexibility in learning novel categories in the open-vocabulary object detection. Thus, we opt for a continuously updating mechanism that could enable the teacher model to evolve and adapt to novel concepts, thereby enhancing its effectiveness in a broader detection scenarios.
Exponential Moving Average. Inspired by the literature [tarvainen2017mean, he2020momentum], we exploit a momentum update strategy for optimizing teacher model, that is, the teacher model is updated by an exponential moving average of the student model during training. Formally, denoting the parameters of the teacher model as $\theta^{\prime}$ and those of the student model as $\theta$. Only the parameters $\theta$ are updated via back-propagation, and parameters $\theta^{\prime}$are updated as a weighted average of successive weights of $\theta$ at each training iteration $t$:
$\theta_{t}^{\prime}=\alpha\theta_{t-1}^{\prime}+(1-\alpha)\theta_{t},$ (3)
where $\alpha\in[0,1)$ is a momentum coefficient.
The momentum update described in Equ. (3) enhances the teacher model in several significant ways compared to a frozen teacher model: 1) it accumulates historical information of the student model without additional training costs, thereby obtaining better intermediate representations and more robust predictions [tarvainen2017mean, liu2021unbiased]; 2) it enables the evolving teacher model to leverage unlabeled data effectively, enhancing the performance of the model with fewer annotated images; and 3) it facilitates on-line learning, allowing the model to continuously adapt and scale to an increasing number of novel concepts.
Consistency Training. Following [softteacher_xu2021end, sohn2020fixmatch], we utilize different augmentation strategies for pseudo labeling. As shown in Fig. 5, weak augmentation involves only random flip, as highlighted by the orange box in the image. On the other hand, strong augmentation incorporates a wide range of augmentation techniques, including both geometric and color space augmentations such as Random Rotate, Auto Contrast, and others.
We adopt Consistency training [jeong2019consistency, sohn2020simple] to guide the student model to make the same prediction as the teacher model makes. Specifically, given an unlabeled image $\mathbf{x}$, it first goes through weak augmentation and serves as the input for the localization teacher model to generate pseudo-labels. The pseudo-labels are then used to supervise the student model, which takes the same image with a strong augmentation as input. Formally, the teacher model’s prediction is defined as $\mathbf{\hat{y}}=f(T^{\prime}(\mathbf{x}),\theta^{\prime})$. The objective is to minimize the consistency cost $J$ between $\mathbf{\hat{y}}$ and the predictions of the student model, i.e.,
@@ 3.2.2 Localization Teacher
where $T$ and $T^{\prime}$ denote strong and weak augmentation, respectively. The cost function $J$ is optimized using stochastic gradient descent (SGD). A detailed explanation of the loss function is provided in Sec. 3.3.2.
3.2.3 Dynamic pseudo-label Queue
@@
Figure 6: Workflow of the dynamic label queue. Step1: filter certain high-quality proposal boxes generated by the localization teacher model, and employ the RemoteCLIP to classify corresponding cropped images as pseudo-labels. Step2: dynamically update the pseudo-labels into the queue, and randomly sample a batch of pseudo-labels for the student model training.
@@
The dynamic label queue is devised for storing, dynamically updating, and transferring pseudo-labels. At each training iteration, the localization teacher model and external teacher model generate high-quality pseudo-labels for the unlabeled images, which are then pushed into the label queue. Following this, a random batch of pseudo-labels is sampled from the dynamic queue for the student model training, as illustrated in Fig. 6.
@@
Generate pseudo-labels. The localization teacher model takes unlabeled images with weak augmentation $T^{\prime}(\mathbf{x})$ as input, where the backbone first extracts features, the RPN then generates a set of proposals $\{p_{i}\}$ and extracts RoI features $\mathbf{f}^{P}=\Phi_{\text{RPN}}\circ\Phi_{\text{ENC}}(T^{\prime}(\mathbf{x}))$ of these proposals through RoI pooling. Subsequently, the class-agnostic regression branch predicts the refined coordinates $\{\hat{b}_{i}\}$. To improve the accuracy of pseudo-labels while reducing computational complexity, we use a box filter to select $k$ candidates and obtain a set of image crops $\{\mathbf{x}^{I}_{i}\}_{i=1}^{k}$ by cropping the corresponding regions of the image. The category labels for these regions are then predicted by the RemoteCLIP.
@@
For each cropped image $\mathbf{x}^{I}_{i}$, the visual feature $v_{i}$ is computed via the visual encoder of the RemoteCLIP, denoted as $v_{i}=\Phi_{\text{clip}}^{\text{visual}}(\mathbf{x}^{I}_{i})$. At the same time, text embeddings $t_{j}$ corresponding to category descriptions are generated, as detailed in Sec. 3.2. The prediction probability for each category $j$ given the cropped image $i$ is the softmax value for similarity between the visual feature and text embeddings:
@@
We further apply a threshold $p_{0}$ to filter pseudo-labels,
$\hat{y}$ $=\left\{(\hat{b}_{i},\hat{c}_{i})\mid\hat{p}_{i}\geq p_{0}\right\}_{i=1}^{\hat{k}}$, (6)
$\hat{c}_{i}$ $=\arg\max_{j}\hat{p}_{ij},\ \hat{p}_{i}=\max_{j}\hat{p}_{ij},$
[…]
@@
This dynamic process allows the queue to progressively accumulate richer and more accurate pseudo-labels as the model iterates, with visualization results shown in Fig. 12. The continuous improvement benefits the entire system, enabling the student and localization teacher models to learn from diverse, high-quality pseudo-labels.
3.3 Horizontal Open-vocabulary Detection
Horizontal object detection aims to detect all objects in a image and represent their location with horizontal bounding boxes, i.e., $b_{i}\in\mathbb{R}^{4}$. To enable open-vocabulary detection, we build upon the Faster R-CNN framework [ren2015faster] and incorporate external knowledge from unlabeled images, which may contain information about novel categories. Specifically, the localization teacher model first generates a set of box candidates, which are then filtered based on their confidence scores. Only those with high confidence are forwarded to the external teacher model for category prediction. The model is then optimized using a hybrid training mechanism that leverages both labeled data and unlabeled data with these high-quality pseudo-labels.
3.3.1 Horizontal Box Selection Strategies
[…]